# Quantum Causality: Resolving Simpson’s Paradox with $\mathcal{DO}$-Calculus

### v1

@@ III-C Result 2: Scalability on a Complex Network (Simulation)
Figure 4: Quantifying confounding bias in the 10-qubit healthcare simulation. This figure compares the treatment effect ($\Delta P$) calculated from raw observational data against results from conventional stratification and our quantum causal intervention. The observational data (‘Overall (Obs.)’) significantly underestimates the true causal effect, yielding an effect of +0.377, whereas the true effect revealed by the intervention (‘Causal Intervention’) is +0.486. Simple stratification by a single variable provides only a partial and inconsistent correction; stratifying by Region yields an effect of +0.406, while stratifying by the stronger confounder, Age, yields +0.497. This demonstrates the robustness of the quantum $\mathcal{DO}$-calculus approach in isolating the true causal effect from complex confounding structures. The horizontal dashed line indicates the true causal effect for reference. Error bars represent 95% confidence intervals calculated over 10 trials.
@@
IV Discussions
In this work, we have demonstrated a quantum algorithmic framework for causal inference, validating its ability to resolve statistical paradoxes and quantify confounding bias through both high-fidelity simulations and a proof-of-principle experiment on a quantum processor. Beyond the specific numerical outcomes, our findings support a deeper interpretation of quantum circuits as direct physical analogues for SCMs. As established in our foundational work [5], this framework uses superposition to represent the probabilistic nature of causal variables. Crucially, both the arrows representing true causal effects and those forming the confounding structures that induce spurious correlations are realized by the same physical mechanism: entanglement generated via controlled gates. From this perspective, the $\mathcal{DO}$-operator is implemented via the circuit surgery procedure—a structural modification that creates a new physical system corresponding to the interventional world. Building on this interpretation, the following sections will analyze the specific biases observed in our models, position our work relative to classical methods and its implications for AI, and discuss the limitations and future directions of this approach.
IV-A Interpretation of Confounding Bias in Causal Models
@@
IV-B Implications and Comparison with Classical Methods
@@
The ability to perform physical causal interventions on a quantum computational substrate has significant implications across several scientific and technological domains. In machine intelligence, our framework offers a new tool for enhancing explainability (XAI) and algorithmic fairness. By moving beyond correlational models, it provides a means to probe the true causal drivers behind a model’s decisions, helping to identify and mitigate biases learned from confounded data. Beyond AI, this approach could be applied to complex network problems in fields such as systems biology, materials science, and econometrics, where simulating interventions to understand the effects of a specific gene, material property, or policy is a central challenge.
@@
This computational approach to intervention stands in contrast to the gold standard of classical experimental design: the RCT. While RCTs achieve causal identification by physically randomizing subjects, they can be prohibitively expensive, time-consuming, unethical (e.g., forcing subjects to adopt harmful behaviors), or physically impossible (e.g., re-running historical events). Our quantum framework, acting as a “computational laboratory,” allows such interventions to be simulated efficiently on a validated causal model. This enables researchers to explore a vast range of “what-if” scenarios that are inaccessible to traditional experimental methods.
@@
It is important to position our quantum approach relative to established classical methods for causal inference, such as covariate adjustment based on the back-door criterion or instrumental variable analysis [4]. For systems where the causal graph is known and the number of variables is small enough to be handled by classical computers, these statistical methods are both efficient and sufficient. However, the advantage of a quantum approach emerges in scenarios where the system’s complexity becomes a prohibitive barrier for classical computation. For causal networks with a vast number of variables and highly intricate dependencies, the state space can grow exponentially, making it intractable to classically represent the joint probability distribution or sample from it. By leveraging quantum properties like superposition and entanglement, a quantum processor can potentially explore these high-dimensional causal structures more efficiently. While the current work does not demonstrate a quantum advantage, our successful experimental validation serves as a foundational proof-of-principle, establishing a pathway toward a potential “quantum advantage for causal inference” on future, fault-tolerant quantum hardware.
IV-C Limitations and Future Directions
@@
While our work establishes a viable framework for quantum causal inference, it is important to clarify its scope and current limitations. The experiments presented in this paper focus on causal effect estimation and therefore presuppose that the underlying causal graph is known a priori, which is a standard assumption for applying the $\mathcal{DO}$-calculus. However, this does not mean the framework is inapplicable when the true graph is unknown. Indeed, the ability of our circuit-based approach to act as a “computational laboratory” allows it to serve as a powerful tool for causal model testing. Researchers can formulate several competing causal hypotheses, compile each into a distinct quantum circuit, and compare the simulated observational data from each circuit against real-world data to determine the most plausible causal model. While a full exploration of such a quantum-enhanced hypothesis testing workflow is beyond the scope of this paper, it represents a significant direction for future research.
@@
Second, while we have provided a successful proof-of-principle on current quantum hardware, a significant gap remains between the performance on small-scale problems and the requirements for achieving a practical quantum advantage on large-scale problems with noisy intermediate-scale quantum (NISQ) devices.
@@
These limitations directly highlight several exciting avenues for future research. A primary direction is the development of quantum algorithms for causal discovery, which would aim to learn the causal graph itself from quantum observational and interventional data, thereby removing our framework’s main prerequisite. Another critical research path is the design of noise-robust causal inference circuits. This could involve co-designing algorithms with specific hardware properties in mind or integrating quantum error mitigation and correction techniques to improve the fidelity of results on current and future NISQ processors. Progress in these areas will be crucial for moving quantum causal inference from a proof-of-principle, as demonstrated here, to a powerful tool for scientific discovery in complex systems.
V Conclusion
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (Ministry of Science and ICT (MSIT)), grant number 2020R1F1A1067619. This research was supported by ‘Quantum Information Science R&D Ecosystem Creation’ through NRF funded by the Korea government (MSIT), grant number 2020M3H3A1110365.
Appendix
Quantum Gate Parameters for the Causal Models

### v2

@@ IV-C Result 2: Scalability on a Complex Network (Simulation)
Figure 4: Quantifying confounding bias in the 10-qubit healthcare simulation. This figure compares the treatment effect ($\Delta P$) calculated from raw observational data against results from conventional stratification and our quantum causal intervention. The observational data (‘Overall (Obs.)’) significantly underestimates the true causal effect, yielding an effect of +0.377, whereas the true effect revealed by the intervention (‘Causal Intervention’) is +0.486. Simple stratification by a single variable provides only a partial and inconsistent correction; stratifying by Region yields an effect of +0.406, while stratifying by the stronger confounder, Age, yields +0.497.
@@
For full reproducibility, the specific gate parameters (rotation angles $\theta$) used to construct both the 3-qubit and 10-qubit causal models, along with the compilation recipe used to derive them from conditional probabilities, are detailed in Appendix A and F.
V Discussions
In this work, we have demonstrated a quantum algorithmic framework for causal inference, validating its ability to reproduce the classical resolution of Simpson-type paradoxes and quantify confounding bias through both high-fidelity simulations and a small-scale experimental demonstration on a quantum processor. Beyond the specific numerical outcomes, our findings support a deeper interpretation of quantum circuits as direct physical analogues for SCMs. As established in our foundational work [8], this framework uses superposition to represent the probabilistic nature of causal variables. Crucially, both the arrows representing true causal effects and those forming the confounding structures that induce spurious correlations are realized by the same physical mechanism: entanglement generated via controlled gates. From this perspective, the $\mathcal{DO}$-operator is implemented via the circuit surgery procedure—a structural modification that creates a new physical system corresponding to the interventional world. Building on this interpretation, the following sections will analyze the specific biases observed in our models, position our work relative to classical methods and its implications for AI, and discuss the limitations and future directions of this approach.
V-A Interpretation of Confounding Bias in Causal Models
@@
V-B Generalization to a Compositional Causal Calculus
The results presented for Simpson’s paradox are a specific instance of a more general capability of our framework to act as a compiler for causal reasoning. As detailed in Sec. III-B, each primitive operation of the $\mathcal{DO}$-calculus is assigned a fixed circuit-level semantics: $\mathcal{DO}$-interventions are implemented via circuit surgery on the observational circuit, conditioning on $Z=z$ corresponds to measuring the $Z$-register and classically post-selecting the desired outcome, and marginalization over $Z$ is realized by summing the resulting measurement statistics.
The key observation is that these primitives compose. Complex causal queries such as the backdoor adjustment formula
$P(y\mid\mathcal{DO}(x))\;=\;\sum_{z}P(y\mid x,z)\,P(z)$
used in our Simpson-type example are built from exactly these ingredients. As formally shown in Appendix E, any valid expression obtained from an SCM by combining interventions, conditioning, and marginalization in the sense of the $\mathcal{DO}$-calculus is faithfully reproduced by the corresponding composition of our circuit primitives. This positions our framework not merely as a tool for resolving a single paradox, but as an experimentally validated mapping from abstract causal calculus to executable quantum circuits.
V-C Implications and Comparison with Classical Methods
@@
The ability to perform physical causal interventions on a quantum computational substrate has significant implications across several scientific and technological domains. In machine intelligence, our framework offers a new tool for enhancing explainability and algorithmic fairness. By moving beyond correlational models, it provides a means to probe the true causal drivers behind a model’s decisions, helping to identify and mitigate biases learned from confounded data. Beyond AI, this approach could be applied to complex network problems in fields such as systems biology, materials science, and econometrics, where simulating interventions to understand the effects of a specific gene, material property, or policy is a central challenge.
@@
This computational approach to intervention stands in contrast to the gold standard of classical experimental design: the RCT. While RCTs achieve causal identification by physically randomizing subjects, they can be prohibitively expensive, time-consuming, unethical (e.g., forcing subjects to adopt harmful behaviors), or physically impossible (e.g., re-running historical events). Our quantum framework, acting as a “computational laboratory,” allows such interventions to be simulated on a validated causal model. This enables researchers to explore a vast range of “what-if” scenarios that are inaccessible to traditional experimental methods.
@@
It is important to position our quantum approach relative to established classical methods for causal inference, such as covariate adjustment based on the back-door criterion or instrumental variable analysis [7]. For systems where the causal graph is known and the number of variables is small, these statistical methods are both efficient and sufficient. The motivation for a quantum approach, however, lies in scenarios where the system’s complexity poses a significant challenge to classical methods. For causal networks with a large number of variables and highly intricate dependencies, the state space can grow exponentially, which can make it computationally expensive to represent and sample from the full joint probability distribution classically. A quantum approach, by leveraging properties like superposition and entanglement, offers a fundamentally different paradigm for representing and manipulating these high-dimensional causal structures. In this work we remain agnostic about any potential speedups and focus instead on showing that SCM-style graph surgery and $\mathcal{DO}$-operations can be faithfully implemented and empirically validated on present-day NISQ processors.
V-D Limitations and Future Directions
@@
While our work establishes a viable framework for quantum causal inference, it is important to clarify its scope and current limitations. The experiments presented in this paper focus on causal effect estimation and therefore presuppose that the underlying causal graph is known a priori, which is a standard assumption for applying the $\mathcal{DO}$-calculus. However, this does not mean the framework is inapplicable when the true graph is unknown. Indeed, the ability of our circuit-based approach to act as a computational laboratory allows it to serve as a powerful tool for causal model testing. Researchers can formulate several competing causal hypotheses, compile each into a distinct quantum circuit, and compare the simulated observational data from each circuit against real-world data to determine the most plausible causal model. While a full exploration of such a quantum-enhanced hypothesis testing workflow is beyond the scope of this paper, it represents a significant direction for future research.
@@
Second, while we have provided a successful proof-of-principle on current quantum hardware, a significant gap remains between the performance on small-scale problems and the requirements for achieving a practical quantum advantage on large-scale problems with NISQ devices.
Finally, throughout this work we deliberately restrict attention to block-diagonal circuits that act within a classical sector of the Hilbert space; relaxing this constraint to allow non-block-diagonal dynamics with genuinely quantum confounders is a natural direction for future work and would clarify in which regimes quantum resources can offer benefits beyond classically simulable SCMs.
@@
Taken together, these limitations highlight several concrete avenues for future research. A primary direction is the development of quantum algorithms for causal discovery, which would aim to learn the causal graph itself from quantum observational and interventional data, thereby removing our framework’s main prerequisite. Another critical research path is the design of noise-robust causal inference circuits. This could involve co-designing algorithms with specific hardware properties in mind or integrating quantum error mitigation and correction techniques to improve the fidelity of results on current and future NISQ processors. Progress in these areas will be crucial for moving quantum causal inference from a small-scale experimental validation, as demonstrated here, to a powerful tool for scientific discovery in complex systems.
VI Conclusion
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (Ministry of Science and ICT (MSIT)), grant number 2020R1F1A1067619. This research was supported by ‘Quantum Information Science R&D Ecosystem Creation’ through NRF funded by the Korea government (MSIT), grant number 2020M3H3A1110365. The authors acknowledge Google’s Gemini for its assistance during the preparation of this manuscript. The AI model was utilized for refining the language and structure of several paragraphs, and improving Python code for circuit visualization.
Appendix A Quantum Gate Parameters for the Causal Models
@@
Appendix B Theoretical Framework and Notation
In this appendix, we provide the formal definitions and theoretical guarantees that underpin the quantum causal inference framework used in the main text. We begin by establishing the core notation.
Setting and Notation. Let $\mathcal{M}=(G,\{P(X_{i}\mid\mathrm{Pa}(X_{i}))\}_{i=1}^{n})$ be a binary structural causal model (SCM) with a DAG $G$ over variables $X_{1},\dots,X_{n}$ and parents $\mathrm{Pa}(X_{i})$. We compile $\mathcal{M}$ into a quantum circuit $\mathcal{C}(G,\Theta)$ on $n$ qubits by assigning one qubit per variable and, for each $i$ in a topological order of $G$, appending a (multi-)controlled rotation that realizes the local conditional:
$\sin^{2}\!\big(\tfrac{\theta_{i}(pa)}{2}\big)$ $=P(X_{i}{=}1\mid\mathrm{Pa}(X_{i}){=}pa),$
$U_{i}(pa)$ $:=\mathrm{CRY}_{i}\!\big(\theta_{i}(pa)\big),$
[…]
Appendix C Operator–Theoretic Embedding (Classical Diagonal Sector)
Definition (Embedding). For a finite binary SCM on DAG $G$, assign one qubit per variable with computational basis $\{\lvert 0\rangle,\lvert 1\rangle\}$. For each arrow $A\!\to\!B$, implement a controlled rotation on $B$ that is active only on a specified control pattern of $A$ (and other parents if present). Readout uses computational–basis projectors.
Lemma (Diagonal/classical form). Under this embedding, all instruments are block–diagonal in the computational basis. Consequently, the global state remains a classical mixture over basis states and $\operatorname{Born}(\rho,E)=\mathrm{Tr}(\rho E)$ reduces to classical sums; i.e., Born probabilities coincide with classical SCM marginals.
Proof sketch. Controlled rotations conditioned on disjoint classical patterns act on orthogonal subspaces; no off–diagonal coherences are created across patterns. Composition preserves block–diagonality, hence Born probabilities equal classical marginals. $\square$
Appendix D Correctness of Circuit Surgery: Proposition and Proof Sketch
Proposition 1 (Correctness of circuit surgery for the compiled family).
For the circuit $\mathcal{C}(G,\Theta)$ constructed above, the following hold.
  1. 1.
[…]
    $\Pr_{\mathcal{C}}\big[X{=}x\big]=\prod_{i=1}^{n}P\big(x_{i}\mid x_{\mathrm{Pa}(X_{i})}\big)=:P_{\mathcal{M}}(x).$
  2. 2.
[…]
    $\Pr_{\mathcal{C}^{\mathcal{DO}}}\big[X{=}x\big]$ $=\mathbf{1}[x_{S}{=}s]\prod_{i\notin S}P\big(x_{i}\mid x_{\mathrm{Pa}(X_{i})\setminus S},\,s\big)$
    $=:P_{\mathcal{M}}\big(x\,\big|\,\mathcal{DO}(S{=}s)\big).$
Proof sketch.
We prove (1) and (2) by induction over a topological order of $G$.
[…]
(2) Interventional (circuit-surgery) case. Deleting all incoming controls to $S$ and fixing $X_{S}{=}s$ precisely implements the truncated factorization: the factors $P(X_{j}\mid\mathrm{Pa}(X_{j}))$ for $j\in S$ are replaced by point masses $\mathbf{1}[x_{j}{=}s_{j}]$, and all descendants retain their conditionals with $S$ clamped to $s$. By the same inductive argument as above, each downstream controlled rotation now conditions on $(x_{\mathrm{Pa}(X_{i})\setminus S},s)$, yielding the product $\mathbf{1}[x_{S}{=}s]\prod_{i\notin S}P(x_{i}\mid x_{\mathrm{Pa}(X_{i})\setminus S},s)$, which equals $P_{\mathcal{M}}(x\mid\mathcal{DO}(S{=}s))$ by the standard truncated-factorization semantics of interventions. ∎
D-A Remark: Scope and Provenance
The intervention operator realized as circuit surgery follows our companion work on Entanglement as Super-Confounding [8], where the operator is introduced and motivated conceptually. Here we instantiate it within an SCM-to-circuit compiler and provide the above correctness guarantee for the compiled circuit family used in the main text. We focus on functional correctness of compiled $\mathcal{DO}$-semantics on hardware; a detailed noise analysis is beyond the present scope.
Appendix E Compositional Semantics for $\mathcal{DO}$, Conditioning, and Marginalization
[…]
$P^{\mathrm{circ}}(\cdot)\;=\;P^{\mathrm{SCM}}(\cdot).$
Proof sketch. Item (i) is Proposition 1(2). For (ii), post–selection implements restriction to the Z=zZ{=}z block (block–diagonal lemma), yielding $P(\cdot\mid Z{=}z)$. For (iii), linearity of trace gives convex mixing over measured outcomes, which equals classical summation. Closure under sequential/parallel composition yields the claim. $\square$
Corollary (Backdoor adjustment via compiled circuit). For the Simpson DAG with confounder $Z$,
$P^{\mathrm{circ}}(y\mid\mathcal{DO}(x))$ $=\sum_{z}P(y\mid x,z)\,P(z)$
$=P^{\mathrm{SCM}}(y\mid\mathcal{DO}(x)).$
Proof. Measure $Z$, evaluate $P(y\mid x,z)$ on each branch, and classically mix with weights $P(z)$. Apply the theorem. $\square$
Appendix F Compilation Recipe: From Conditionals to $\mathrm{CRY}$ Angles
The tables in the main appendix list the concrete angles actually used. For completeness, we record the generic mapping used by our compiler.
Single–parent arrow $A\!\to\!B$.
Initialize $B$ in $\lvert 0\rangle$. If $P(B{=}1\mid A{=}1)=p_{1}$, apply a single $\mathrm{CRY}_{A\to B}(\theta_{1})$ controlled on A=1A{=}1 with
$\theta_{1}\;=\;2\arcsin\!\sqrt{p_{1}},\qquad\text{so that}\quad\sin^{2}\!\bigl(\tfrac{\theta_{1}}{2}\bigr)=p_{1}.$
If you need a link conditioned on A=0A{=}0, wrap control by $X$ on $A$: $\mathrm{CRY}XX$.
Two–parent arrow $(A,B)\!\to\!C$.
For each parent pattern $(a,b)\in\{0,1\}^{2}$, if the desired conditional is $P(C{=}1\mid a,b)=p_{ab}$, apply a mutually exclusive doubly–controlled rotation $\mathrm{CRY}_{(a,b)\to C}(\theta_{ab})$ that fires only on (A=a,B=b)(A{=}a,B{=}b), with
$\theta_{ab}\;=\;2\arcsin\!\sqrt{p_{ab}}.$
Because the controls are disjoint, only one gate can be active on any branch; the order of these gates is therefore immaterial, and the resulting probability on that branch equals $p_{ab}$.
Multi–parent generalization.
For $k$ parents, decompose into $2^{k}$ mutually exclusive control patterns and set $\theta_{\mathbf{a}}=2\arcsin\sqrt{p(C{=}1\mid\mathbf{a})}$ for each pattern $\mathbf{a}$. Controls on logical 0 are implemented by XX–wrapping the corresponding qubit.
Remark. This “exclusive–control” construction is the one instantiated by the angle tables. When hardware connectivity requires it, standard decompositions (e.g., ancilla–assisted Toffoli ladders) are used without altering the angles; exclusivity of controls is preserved by construction.